### v1

Abstract
Code cloning is frequently observed in software development, often leading to a variety of maintenance and security issues. While substantial research has been conducted on code cloning in traditional software, to the best of my knowledge, there is a lack of studies on cloning in VR software that consider its unique nature, particularly the presence of numerous serialized files in conjunction with the source code. In this paper, we conduct the first large-scale quantitative empirical analysis of software clones in 345 open-source VR projects, using the NiCad detector for source code clone detection and large language models (LLMs) for identifying serialized file clones. Our study leads to a number of insights into cloning phenomena in VR software, guided by seven carefully formulated research questions. These findings, along with their implications, are anticipated to provide useful guidance for both researchers and software developers within the VR field.
keywords:
Virtual Reality (VR), Code Clone, Large Language Model
1 Introduction
In recent years, Virtual Reality (VR) has gained substantial traction, finding significant applications across diverse domains such as gaming, healthcare, education, and entertainment[1, 2, 3]. An increasing number of applications are being created and distributed on leading app platforms (e.g., Google Play, Apple Store, Oculus), collectively reaching approximately 200 million downloads worldwide [4]. However, research on the quality of VR software remains limited, with code clone—a common yet often overlooked practice in software development—being particularly unexplored within this context.
Code clone generally refers to the repetition of identical or similar code snippets in multiple places within a software project, which sometimes may pose critical challenges for software security[5], maintainability[6] and changeability[7]. While there has been extensive research on the issue of code cloning in traditional software[8, 9, 10, 11, 12, 13, 14, 15, 16], to the best of my knowledge, our prior work[17] is the first quantitative empirical study of code cloning in open-source VR software. Due to space limitations, the study [17] limits its scope to source-code level code cloning. Nonetheless, VR software exhibits a unique characteristic compared to traditional software, that is, it incorporates a large number of serialized files, which are pivotal for data representation and storage in VR projects. Cloning of these files may also introduce various issues and potential risks. For instance, in cases where identical small balls are created by copying the same .prefab, any updates (e.g., to size or material) must be applied to each copy manually, leading to potential inconsistencies. Additionally, flaws in the original .prefab, like missing colliders or script errors, will affect all copies, making debugging more difficult. This reinforces the need for clone detection in such non-traditional code files in tandem with source code, a topic that is still surprisingly under-researched.
In this study, we conduct the first quantitative empirical analysis of code cloning in 345 open-source VR projects, conducting clone detection in both source code and serialized files. For the source code portion of VR software, we use the NiCad tool [18, 15] for analysis, while for serialization files— which are difficult to detect using traditional methods— we utilize Large Language Models (LLMs) for analysis. Through seven carefully formulated research questions (RQs) and a set of proposed metrics, our study uncovers a number of insights into cloning phenomena in VR software and highlights the unique challenges of VR software development, offering practical guidance for researchers, developers, and maintainers within the VR field. Some of our insights are highlighted as follows: (i) Both source code cloning and serialized file cloning in VR software show a consistent pattern, with larger projects generally exhibiting more clones. (ii) The main difference between code clones in VR software and regular software lies in the presence of serialized files, which significantly reduce the level of source code cloning in a project. (iii) C# is the most commonly used programming language in open-source VR projects, whereas C is more likely to cause source code cloning issues. (iv) Third-party libraries introduced in VR applications often contribute to source code cloning issues, sometimes constituting the majority of the cloning results. (v) Inter-version code cloning is more common than intra-version code cloning, introducing significant security risks due to the potential propagation of vulnerabilities across multiple versions of the software. (vi) The cloning behavior of asset files varies with complexity. Complex files like scenes have lower clone levels, simpler ones like materials have higher, and prefab files fluctuate with development style.
To sum up, this paper provides the following main contributions:
  • We highlight the unique nature of code cloning in VR software relative to traditional software and conduct a thorough quantitative empirical analysis of both source code clones and serialized asset file clones in our constructed dataset of 345 open-source VR projects.
  • We formulate a suite of metrics to quantify code clones from various angles, enabling a comprehensive quantitative evaluation when used together.
  • We propose a detection method based on large language models to identify clones in the unique serialized asset files of VR software, presenting a novel approach to code clone detection in this domain.
  • We derive several insights from the experimental results that deepen our understanding of VR software, providing practical guidance for researchers, developers, and maintainers in the VR domain.
Paper Organization. The paper is structured as follows. Section 2 provides the background of VR software. Section 3 explains our methodology. Section 4 presents the results of seven research questions. Section 5 suggests future work. Section 6 discusses related work. Section 7 concludes the paper.
2 Background
In this section, we provide a brief introduction to VR software and discuss methods for detecting code cloning issues within it.
2.1 VR Software Architecture
VR software tightly integrates source code with serialized asset files, collectively shaping the application’s functionality and immersive features. During runtime, source code engages directly with assets to perform essential tasks, such as loading and rendering 3D models, triggering and managing animations, and processing user inputs to deliver interactive experiences. In the realm of VR software development, platforms like Unity [19, 20, 21] play a pivotal role, offering a comprehensive architectural model that synthesizes code-directed operational behaviors with immersive, asset-structured environmental elements.
Source Code. In VR software, the source code primarily consists of scripts that define the core functionality of the application. Typically written in C#, these scripts govern object behavior, manage user interactions, and implement key features such as rendering, physics simulations, and input handling. In Unity, source code adopts a content-logic decoupling and component-based architecture, where individual scripts are assigned to game objects to regulate specific aspects of their functionality. While this design enhances reusability, it can also result in redundancy, particularly in large-scale or collaborative development contexts.
Serialized Asset Files. In VR software, asset files serve as the backbone for delivering visual, auditory, and interactive content. In Unity, assets are serialized for efficient storage and retrieval. Common serialized files in VR software include asset files such as material files (.mat, .shader, .cginc), scene files (.unity, .scene, .navmesh), resource files (.fbx, .wav, .png), and template files (.prefab), along with configuration files (.json, .ini) and metadata files (.meta) bound one-to-one with asset files [19]. While the serialized structure ensures platform interoperability and smooth resource integration, its component-based design can result in redundancy across projects or scenes.
2.2 VR Code Clone
In light of the aforementioned VR software architecture, this paper extends the concept of code clone beyond its conventional scope at the source code level to include the cloning of data embedded in serialized asset files.
2.2.1 Source Code Clone
In the context of software development, source code clone denotes the duplication of similar code fragments within a codebase, where code fragment refers to a continuous segment of source code. Generally, there are four clone forms[22, 23, 14, 24, 25]: (i) identical clone, referring to identical code fragments except for variations in comments and layout; (ii)lexical clone, referring to changes in identifier names and lexical values on the basis of the identical clone; (iii)syntactic clone, referring to syntactically similar code fragments that add statements of mutual addition, modification, or deletion on the basis of the identical and lexical clone; (iv)semantic clone, referring to syntactically different code fragments that implement the same functions. Note that the first three clone types indicate textual similarity whereas the last type reflects functional similarity.
Various approaches for source code clone detection have been proposed in the literature, as elaborated below. (i)Textual-based clone detection [26] compares the raw text of code to identify clones, typically relying on string-matching techniques. This approach is suitable for detecting explicit and simple clones, such as exact copy-pasted code. Representative implementations of this method include Simian [27] and Duplo [28]. (ii)Token-based clone detection [29, 30] transforms the source code into tokens (lexical units) before comparison. By ignoring non-semantic elements such as whitespace, comments, and optionally identifiers, it identifies clones based on token sequences. Representative implementations of this method include NiCad [18, 15] and SourcererCC [25]. (iii)Syntax-based [31, 32] clone detection analyzes the abstract syntax tree (AST) of source code to compare the syntactic structure of code fragments, allowing it to eliminate formatting differences and detect syntactically similar code. Representative implementations of this method include CloneDR [33, 34] and Deckard [35]. (iv)Semantic-based clone detection [36, 37] identifies clones by analyzing the semantic structure of code fragments, focusing on whether the code fragments are functionally equivalent rather than merely syntactically similar. Representative implementations of this method include SCAM [38]. (v)Learning-based clone detection [39, 32] leverages machine learning or deep learning techniques to train models that automatically identify code clones. Recently, a rising number of studies have explored the use of Large Language Models (LLMs) for detecting code clones. [40]. Representative implementations of this method include CodeBERT [41, 42], BigCloneBench [43] and CodeGPTSensor [44].
It is worth noting that the first four detection methods are capable of detecting identical clones, lexical clones, and certain types of syntactic clones. In contrast, learning-based methods, including those utilizing LLMs, excel at identifying partial semantic clones but are less effective in detecting the other three types of clones compared to traditional approaches[45, 40, 46, 47]. Considering these facts, this study intends to employ Nicard, the state-of-the-art technique within traditional methods, for source code clone detection.
2.2.2 Serialized Asset Files Clone
Recall that serialized asset files in VR software serve as data containers, systematically defining the core visual and interactive aspects of a virtual environment. The structured nature of asset files is prone to cause “data clone”, where configurations, asset references, or component settings are duplicated across files. Therefore, serialized asset file clone focuses on identifying duplicated or near-duplicated asset data, such as repeated material definitions, component hierarchies, or configuration settings. Existing detection techniques used for identifying data clone include: (i)Content-based methods, which typically construct feature vectors of the content for similarity comparison [48, 49, 50, 51]. Representative implementations of this method include Scikit-learn [52]. (ii)Fingerprint-based methods, such as rolling hashes [53], which maps similar content to an identical hash value, and MD5 [54], which detects file clones by comparing the similarity of hash values. Representative implementations of this method include LSH [53] and Duplicate File Finder [55]. (iii)Deep learning-based methods, such as word vector models [56, 57, 58] and Transformer-based models [59], which train neural networks to identify textual similarity. Representative implementations of this method include Hugging Face Transformers [60].
Among these methods, content-based methods are simple and easy to implement, but they require significant computational resources and time, limiting their applicability to small-scale texts. Fingerprint-based methods, while computationally efficient, are sensitive to local changes and noise in the text. Deep learning-based methods, on the other hand, require substantial data and time for training, but with the advent of large models, this limitation has been significantly alleviated. In most scenarios, deep learning-based methods outperform the other approaches. The latest studies indicate that GPT-4, owing to its powerful contextual understanding and generative capabilities, often surpasses traditional embedding-based methods and other generative models in text similarity detection tasks, particularly those requiring reasoning and handling contextual variations [61, 62]. Given these facts, we propose using GPT-4o to identify clones in serialized asset files.
3 Methodology
The research methodology overview is shown in Figure 1. We begin by searching for VR projects in GitHub and existing datasets. After selecting 658 projects, we filter them by programming language, reducing the dataset to 416 projects. We then manually inspect and deduplicate these projects, resulting in a final dataset of 345 unique projects. Next, we formulate a set of metrics for measuring VR software cloning from various perspectives. We then propose detection methods for both source code cloning and serialized asset file cloning. Afterward, we conduct code cloning analysis from three angles to answer seven carefully formulated research questions.
Figure 1: Overview of Our Empirical Study.
3.1 Dataset Construction
Despite the availability of several datasets [63, 64, 65, 17] from prior studies, we have reconstructed a dataset tailored to meet four key criteria crucial for the objectives of this study: coverage to include a wide range of VR software projects, diversity to encompass varied types of VR software projects, popularity to focus on prominent GitHub repositories, and usability to enable efficient detection. The generation of the dataset follows four phases, as described below.
Project selection. Given the dynamic accessibility of open-source projects on GitHub, we construct our dataset by combining the datasets disclosed in two latest studies [63, 17] and augmenting them with the latest projects published afterward. The dataset provided by [63] includes 314 open-source VR software projects, comprising 164 independent projects, 63 organizational projects, and 83 academic projects, with 4 of them no longer accessible. The dataset reported by [17] initially consists of 326 popular open-source VR software projects prior to any further processing. We further apply the same filtering criteria as outlined in [66] to incorporate the most recent data, resulting in the inclusion of 22 additional projects to enhance the dataset. By integrating the results of the three methods, we compile a preliminary dataset consisting of 658 open-source VR software projects.
Language filtration. Considering the wide range of development environments for open-source VR projects on GitHub, we propose to identify the projects that can be analyzed by off-the-shelf clone detection tools. For this purpose, we consider NiCad [15], a publicly available, state-of-the-art tool for source code clone detection, supporting languages such as C, C#, Java, Python, PHP, Ruby, Swift, ATL, and WSDL. By applying this tool to the previously obtained dataset, we exclude projects implemented in languages not supported by NiCad, resulting in 416 projects suitable for further analysis.
Manual inspection and deduplication. To ensure dataset accuracy, we manually review the tags and “README.md” files of each project, filtering out non-VR applications. Subsequently, we eliminate duplicate entries within the 416 projects, producing a final dataset of 345 projects, referred to as the VR-345 dataset. The dataset comprises 233 applications and 112 development tools.
3.2 Code Clone Detection
In this section, we establish a set of metrics to quantify the degree of cloning in both source code and serialized asset files, and provide detailed approaches for detecting these two cloning targets.
3.2.1 Metrics for Measuring Code Clone
For source code clone, we focus on code clone at the function granularity. Let $F$ denote the universe of functions in a VR software, and let ${\sf Sim}(func_{i},func_{j})(i\neq j)$ denote the similarity between function $func_{i}\in F$ and function $func_{j}\in F$. We define the concepts of clone function and clone class below.
Clone function. A function $func_{i}\in F$ is deemed a clone function if there exists at least one other function $func_{j}\in F(i\neq j)$ such that the similarity between $func_{i}$ and $func_{j}$ exceeds a predefined similarity threshold $\tau$. Mathematically, a function $func_{i}\in F$ is a clone function under the following condition:
$\exists func_{j}\in F(i\neq j),{\sf Sim}(func_{i},func_{j})>\tau$ (1)
Clone class. A clone class is defined as a set of two or more functions where the similarity between every pair of functions exceeds a predefined similarity threshold $\tau$. Mathematically, a set $C\subseteq F$ ($|C|\geq 2$) is a clone class under the following condition:
$\forall func_{i},func_{j}\in C(i\neq j),{\sf Sim}(func_{i},func_{j})>\tau$ (2)
Grounded in the above definitions, we put forward four metrics to evaluate the degree of source code cloning from both absolute and relative perspectives. The metrics applied to assess the absolute degree of code cloning are as follows:
  • Number of Clone Functions ($\sf NCF$): This metric measures the total number of clone functions in the source code of a VR software, namely
    ${\sf NCF}=|\{func_{i}\in F:\exists func_{j}\in F(i\neq j),{\sf Sim}(f_{i},f_{j})>\tau\}|$ (3)
  • Number of Clone Classes ($\sf NCC$): This metric measures the number of clone classes in the source code of a VR software, namely
    ${\sf NCC}=|\{C_{k}\subseteq F:\forall func_{i},func_{j}\in C_{k}(i\neq j),{\sf Sim}(func_{i},func_{j})>\tau\}|$ (4)
The metrics applied to assess the relative degree of code cloning include:
  • Ratio of Clone Functions ($\sf rcf$): This metric reflects the share of clone functions in a VR software, which is defined as the ratio of the number of clone functions ($\sf NCF$) to the total number of functions ($|F|$) in the software, namley
    ${\sf rcf}={\sf NCF}/|F|$ (5)
  • Ratio of Clone Classes ($\sf rcc$): This metric reflects the share of clone classes in a VR software, which is defined as the ratio of the number of clone classes ($\sf NCC$) to the total number of functions ($|F|$) in the software, namely
    ${\sf rcc}={\sf NCC}/|F|$ (6)
For serialized asset files clone, we focus on data clone at the file granularity. Let $A$ denote the universe of asset files in a VR software, and let ${\sf Sim}^{\prime}(file_{i},file_{j})(i\neq j)$ denote the similarity between asset file $file_{i}\in A$ and asset file $file_{j}\in A$. We define the concepts of clone file and clone group as follows.
Clone file. A serialized asset file $file_{i}\in A$ is deemed a clone file if there exists at least one other asset file $file_{j}\in A(i\neq j)$ such that the similarity between $file_{i}$ and $file_{j}$ exceeds a predefined similarity threshold $\tau^{\prime}$. Mathematically, a serialized asset file $file_{i}\in A$ is a clone file under the following condition:
$\exists file_{j}\in A(i\neq j),{\sf Sim}^{\prime}(file_{i},file_{j})>\tau^{\prime}$ (7)
Clone group. A clone group is defined as a set of two or more serialized asset files where the similarity between every pair of files exceeds a predefined similarity threshold $\tau^{\prime}$. Mathematically, a set $G\subseteq A$ ($|G|\geq 2$) is a clone group under the following condition:
$\forall file_{i},file_{j}\in A(i\neq j),{\sf Sim}^{\prime}(file_{i},file_{j})>\tau^{\prime}$ (8)
Drawing from the definitions outlined above, we introduce a set of metrics to evaluate the extent of serialized asset cloning from both absolute and relative perspectives. The metrics applied to assess the absolute degree of asset cloning are as follows:
  • Number of Clone Assets ($\sf NCA$): This metric measures the total number of clone files in the serialized asset files of a VR software, namely
    ${\sf NCA}=|\{file_{i}\in A:\exists file_{j}\in A(i\neq j),{\sf Sim}^{\prime}(file_{i},file_{j})>\tau^{\prime}|$ (9)
  • Number of Clone Groups ($\sf NCG$): This metric measures the number of clone groups in the serialized asset files of a VR software, namely
    ${\sf NCG}=|\{G_{k}\subseteq A:\forall file_{i},file_{j}\in G_{k}(i\neq j),{\sf Sim}^{\prime}(file_{i},file_{j})>\tau^{\prime}\}|$ (10)
The metrics applied to assess the relative degree of asset cloning include:
  • Ratio of Clone Assets ($\sf rca$): This metric reflects the share of clone files in a VR software, which is defined as the ratio of the number of clone files ($\sf NCA$) to the total number of asset files ($|A|$) in the software, namely
    ${\sf rca}={\sf NCA}/|A|$ (11)
  • Ratio of Clone Groups ($\sf rcg$): This metric reflects the share of clone groups in a VR software, which is defined as the ratio of the number of clone groups ($\sf NCG$) to the total number of asset files ($|A|$) in the software, namely
    ${\sf rcg}={\sf NCG}/|A|$ (12)
  • Clone Index ($\sf CI$): This metric measures the overall similarity of asset files in a VR software, computed by dividing the sum of pairwise similarities by the total number of asset files, namely
${\sf CI}=\frac{\sum_{file_{i},file_{j}\in A,i\neq j}{\sf Sim}^{\prime}(file_{i},file_{j})}{|A|}$ (13)
3.2.2 Source Code Clone Detection
We utilize various parameters of NiCad-6.2 to detect source code clones within the constructed dataset, followed by a detailed evaluation of the detection outcomes to compute relevant clone metrics. Recall that code clones are typically classified into four types: identical, lexical, syntactic, and semantic clones, as described in Section 2.2.1. As shown in Table 1, NiCad-6.2 is capable of detecting the first three types of clones, which are further refined into six distinct subtypes: Type1 represents identical clones; Type2 and Type2c correspond to lexical clones, differing in their handling of identifiers through two methods: blind and consistent; Type3-1, Type3-2, and Type3-2c correspond to syntactic clones, where a dissimilarity threshold, $1-\tau$, with $\tau$ denoting the similarity threshold defined above, is applied to enable the detection of near-miss clones. For example, a dissimilarity threshold of “0.1” allows for up to 10% difference. We set the dissimilarity threshold to “0.3”, which is the default value. The cloning granularity may be set at the function or block level. We opt for function-level granularity to achieve results with greater detail and specificity.
We perform source code clone detection on the VR-345 dataset. The process involves normalizing and parsing source code files for NiCad compatibility, followed by clone analysis to identify clone pairs and classify them into clone classes. For each project, we calculate key metrics, including $\sf NCC$, $\sf NCF$, $\sf rcf$, and $\sf rcc$, to evaluate the clone detection results.
Table 1: NiCad Settings for Source Code Clone Detection
Parameters Target Clone Types
[…]
Dissimilarity Threshold / / / 0.3 0.3 0.3
Identifier Renaming none blind consistent none blind consistent
Granularity function function function function function function
3.2.3 Serialization File Clone Detection.
We employ a two-step Chain of Thought (CoT) [67] technique using the GPT-4o large language model to identify similarities in serialized files for clone detection. Unlike traditional one-step question-answering prompts, this method decomposes complex problems into smaller, sequential steps, where each step’s output informs the next. Studies [68, 69, 70] support the effectiveness of CoT prompting in enhancing reasoning tasks, including clone identification.
Step 1 of the CoT framework requires the large language model to evaluate whether a single file pair qualifies as a clone pair. To facilitate this, four prompt types are developed: simple binary analysis, structure analysis, semantic analysis, and similarity analysis. Details of these prompts are provided in Table 2. Experimental evaluation of these prompts yields several key observations. Simple prompt leads to lower precision, as the model faces challenges in detecting inter-document clones effectively. Structure-based prompt results in high false positive rates, implying that the model is overly responsive to structural similarities within the files. Similarly, semantic dimension prompt also yields high false positive rates, likely due to the simple semantic structure of serialized asset files compared to natural or programming languages, leading to erroneous predictions. In contrast, prompt from the similarity analysis, which enables the model to generate a similarity score for file pairs, demonstrates the best performance and is subsequently chosen as the basis for future detection prompt.
After calculating the similarity among file pairs, we employ Step 2 of the CoT to compute the $\sf NCA$, $\sf NCG$ and $\sf CI$ for each project, as detailed in Table 3. We set the clone threshold ($\tau^{\prime}$) to 80%, following the default recommendation of GPT-4o. The results are subsequently refined through manual inspection and standardization to ensure reliability. Note that the detected serialized files comprise scene files (.unity, .scene, .navmesh), template files (.prefab), material files (.mat, .shader, .cginc), and configuration files (.json, .ini), but exclude resource files (.fbx, .wav, .png) and .meta files. This is because resource files are typically not stored in serialized text format, while .meta files are tightly bound to their corresponding assets.
Table 2: Step 1 of the CoT.
Step Types Prompts
1 Simple Please analyze the file pairs and determine if they are clone files.
[…]
Table 3: Step 2 of the CoT.
Step Metrics Prompts
2 $\sf NCA$, $\sf NCG$ Please analyze the files in the folder and output the number of clone files and clone groups, referencing the cloning information of each file pair from the Output of Step 1.
[…]
3.3 Research Questions Definition
The objective of this study is to systematically investigate the phenomenon of code clones in VR software, aiming to clarify their characteristics, root causes, and implications, and to produce actionable insights for both researchers and practitioners alike. To achieve this goal, we design 7 research questions (RQs) from three perspectives: general cloning in software, cloning specific to source code, and cloning specific to serialized asset files.
[…]
  • RQ1: Which VR projects are subjected to heavy cloning? We begin by examining the impact of different detection methods on clone detection results. Following this, we explore the degree of cloning in different VR software projects, considering both the source code and serialized file levels. Addressing this question will shed light on the projects with the most significant cloning, thereby assisting researchers and developers in prioritizing their efforts.
  • RQ2: What are the main differences in code cloning between VR and non-VR software? We first examine the structural differences between VR and non-VR software, and subsequently investigate how these structural variations contribute to the differences in code cloning between the two types of software. Addressing this question will clarify whether the observed findings are specific to VR software or reflect more universal trends within the software development domain.
  • RQ3: How to identify the most appropriate metrics for VR cloning analysis? We offer an in-depth analysis of the physical interpretations and applicability of the proposed metrics for measuring cloning. By addressing this question, we aim to guide users in selecting the most suitable metrics for various purposes and contexts.
[…]
  • RQ4: Which programming languages dominate VR software development and are more susceptible to code cloning? We analyze the programming languages commonly used in VR software development, the code cloning issues they induce, and the underlying causes, aiming to identify which languages are more prone to cloning and why. Addressing this question will enable us to identify which projects are more prone to code cloning and allow for reflection on how language-level interventions might resolve these cloning challenges.
  • RQ5: How do third-party libraries impact code cloning in VR software? We explore the third-party libraries frequently utilized in VR projects and analyze whether their usage leads to code cloning and the specific patterns of cloning that arise. Addressing this question will help us identify the root causes of code cloning, whether it stems from the use of problematic third-party libraries or from developer practices, thereby allowing us to formulate targeted solutions.
  • RQ6: How do inter-version and intra-version code cloning evolve across different versions of a VR project? We investigate the presence of code cloning both within the same software version and between consecutive versions, exploring the degree of cloning and particularly focusing on the differences in cloning across different granularities of version changes. Addressing this question allows us to understand the progression of code cloning across software versions and uncover hidden security issues within these changes.
Cloning specific to serialized asset files (RQ7) examines the influence of asset types on file cloning.
  • RQ7: How do cloning practices vary across different types of asset files in VR software? We analyze the cloning discrepancies of various types of serialized asset files across multiple distinct projects. Addressing this question provides a finer understanding of asset file cloning, allowing developers to implement tailored maintenance practices for different types of assets.
4 Evaluation
4.1 Overall Analysis
4.1.1 RQ1: Which VR projects are subjected to heavy cloning?
To answer this RQ, we utilize NiCad to detect code clones within the VR-345 dataset and present the most cloned projects from the development tools and applications in Table 4. Specifically, PID 1-10 represent the top ten projects with the highest number of cloned functions in the development tool category, while PID 11-20 are the top ten projects in the VR application category with the most cloned functions. For ease of observation, these projects are arranged in descending order of the total number of functions $N$ within each category.
We begin by examining the impact of different detection methods on the results, revealing a clear and consistent trend in the ${\sf NCF}$ metric (i.e., the number of clone functions): ${\sf NCF}$(Type 3-2) > ${\sf NCF}$(Type 3-2c) > ${\sf NCF}$(Type 3-1), and ${\sf NCF}$(Type 2) > ${\sf NCF}$(Type 2c) > ${\sf NCF}$(Type 1). This suggests that the blind renaming method consistently identifies the most clone fragments, with the consistent renaming method coming in second. This finding exactly aligns with the theoretical expectations outlined in NiCad [15], where blind renaming replaces all identifiers with a generic placeholder “$X_{n}$”, while consistent renaming assigns sequential identifiers “$X_{n}$”($n$ is a sequence number). The rationale is straightforward: if a code fragment is recognized as similar with identifiers renamed as “$X_{n}$”, it will also be recognized as similar when all identifiers are replaced with “$X$”. Additionally, we observe that ${\sf NCF}$(Type 3-x) > ${\sf NCF}$(Type x) for x = 1, 2, 2c. This outcome is intuitive since a fragment meeting a 0% dissimilarity threshold will inherently meet a 30% threshold. Based on these observations, Type 3-2 emerges as the most effective detection method, with its results encompassing those of other types. In terms of the ${\sf NCC}$ metric, there appears to be no clear correlation between the detection types and this metric. As a result, we select Type 3-2 as the standard for code clone detection in the following analyses.
We further observe that the ${\sf NCF}$ values for all twenty projects exceed 100, suggesting that code cloning is a common issue in VR project development. Additionally, the highest number of functions ($|F|$) tend to exhibit the largest volume of code clones (${\sf NCF}$). Specifically, the top three projects with the most functions are “BlenderXR”, “gpac”, and “SoundSpace”, which also show the highest occurrence of code clones. This observation suggests that, generally, the prevalence of code cloning increases as the scale of a project expands, echoing similar trends found in conventional software development.
Table 4: The quantity of clones detected in the source code.
PID Project Name Language Type1 Type2 Type2c Type3-1 Type3-2 Type3-2c $|F|$
[…]
Development Tools
1 BlenderXR C,Python 358 175 1578 601 1486 567 1731 716 4448 1226 3112 889 27822
2 gpac C 6 3 1421 259 1268 220 1195 312 2699 326 2089 278 10170
3 The-Seed-Link-Future C# 44 19 231 84 223 85 201 86 462 166 295 112 8475
4 open-brush C# 38 17 132 60 127 58 183 81 409 163 218 89 7548
5 UnityOculusAndroidVRBrowser C# 36 15 156 53 150 55 107 47 255 86 177 65 3716
6 UnityPlugin C# 12 6 53 21 52 21 60 24 113 43 85 33 2799
7 lovr C 2 1 84 33 77 30 77 31 232 64 148 53 2173
8 UnityGameTemplate C# 20 10 60 25 59 25 53 22 142 47 98 35 1981
9 ViveInputUtility-Unity C# 20 10 62 27 62 27 69 32 125 49 84 36 1863
10 com.xrtk.core C# 3 1 49 18 49 18 42 17 148 51 82 32 1607
Applications
11 SoundSpace C# 2149 1063 2190 1011 2185 1009 2175 985 2225 965 2201 1001 9117
12 RhythmAttack-VR C# 37 17 162 60 149 62 118 48 330 119 221 89 6751
13 Dungeon-VR C# 20 10 100 38 91 35 103 41 254 98 154 61 6470
14 Situated-Empathy-in-VR C# 23 11 140 59 136 58 136 59 388 140 262 100 4788
15 mineRVa C# 18 9 88 35 82 34 53 24 197 71 140 58 4722
16 VR-Escape-Room C# 12 6 53 19 44 16 31 13 120 46 77 30 4067
17 Group6_ProjectNurture C# 27 10 148 49 142 51 102 44 248 82 171 62 3980
18 vrtist C# 51 17 114 42 112 41 141 48 299 103 187 69 3322
19 Terminal C# 19 8 96 27 88 29 71 27 160 49 121 41 2929
20 elite-vr-cockpit C# 0 0 52 18 49 17 30 14 109 39 68 26 2725
To expand our study on cloning in serialized asset files within VR software, we add 10 more VR applications, covering diverse subcategories like games, education, simulators, and AR, ensuring a comprehensive analysis. Development tools are excluded from the analysis of file cloning because they are typically source-code-centric, with asset files serving configuration and descriptive purposes. In contrast, asset files dominate in application projects, with source code primarily existing as scripts to manipulate these assets. Table 5 provides the clone detection results for the 20 selected VR software applications. We observe that the three projects with the highest number of file clones (${\sf NCA}$)—“Dungeon-VR”, “Group6_ProjectNurture” and “OpendagVR2”—are also the ones with the largest number of asset files ($|A|$), following a pattern similar to source code cloning.
Table 5: The quantity of clones detected in serialization files.
PID Project Name $\sf NCA$ $\sf NCG$ $|A|$ $\sf rca$ $\sf rcg$ $\sf CI$
[…]
11 SoundSpace 207 10 15 3 28 391 52.94% 7.16% 11.24%
12 RhythmAttack-VR 286 20 19 5 44 658 43.47% 6.69% 8.17%
13 Dungeon-VR 941 31 20 15 66 1737 54.17% 3.80% 16.32%
14 Situated-Empathy-in-VR 158 1 3 2 6 209 75.60% 2.87% 57.42%
15 mineRVa 235 8 8 3 19 411 57.18% 4.62% 18.67%
16 VR-Escape-Room 154 10 6 4 20 355 43.38% 5.63% 9.69%
17 Group6_ProjectNurture 777 24 15 12 51 1099 70.70% 4.64% 23.28%
18 vrtist 293 21 17 5 43 408 71.81% 10.54% 14.89%
19 Terminal 268 11 13 4 28 450 59.56% 6.22% 23.18%
20 elite-vr-cockpit 97 5 8 1 14 212 45.75% 6.60% 12.35%
21 VRHamsterBall 108 5 10 1 16 194 55.67% 8.25% 16.37%
22 OpendagVR2 699 21 26 10 57 947 73.81% 6.02% 24.05%
23 epicslash 145 7 8 6 21 210 69.05% 10.00% 17.46%
24 CityMatrixAR 70 3 5 2 10 115 60.87% 8.70% 25.13%
25 Procrastination-VR 438 5 5 2 12 478 91.63% 2.51% 69.29%
26 AirAttack 265 8 9 5 22 352 75.28% 6.25% 27.28%
27 GolfVR 56 4 3 2 9 116 48.28% 7.76% 13.84%
28 XR-Keyboard 70 5 2 2 9 159 44.03% 5.66% 15.18%
29 HorrorGame 349 12 27 7 46 506 68.97% 9.09% 13.93%
30 Pokemon-Themed-Kiosk-VR 172 10 4 5 19 239 71.97% 7.95% 22.96%
Insight 1.  Both source code cloning and serialized file cloning in VR software show a consistent pattern, with larger projects generally exhibiting more clones.
4.1.2 RQ2: What are the main differences in code cloning between VR and non-VR software?
To answer this RQ, we suggest exploring the relationship between the quantity of asset files ($|A|$) and the count of clone functions (${\sf NCF}$), as VR software differs from traditional software primarily in its usage of serialized asset files.
As shown in Table 6, the number of cloned functions and serialized asset files for each of the 30 projects is presented. We observe that the development tools Project-1 (“BlenderXR”), Project-2 (“gpac”), and the VR applications Project-11 (“SoundSpace”), Project-24 (“CityMatrixAR”) exhibit significantly higher levels of source code cloning than the rest of the projects. Considering that Project-1 and Project-2 neither use C# nor contain serialized asset files, these two projects can effectively be viewed as resembling non-VR software.
Table 6: The number of asset files and clone functions in the projects.
PID 1 2 3 4 5 6 7 8 9 10
$\sf NCF$ 4448 2699 462 409 255 113 232 142 125 148
$|A|$ 0 0 303 1037 321 222 0 63 93 8
PID 11 12 13 14 15 16 17 18 19 20
$\sf NCF$ 2225 330 254 388 197 120 248 299 160 109
$|A|$ 391 658 1737 209 411 355 1099 408 450 212
PID 21 22 23 24 25 26 27 28 29 30
$\sf NCF$ 74 45 54 196 76 94 46 10 11 2
$|A|$ 194 947 210 115 478 352 116 159 506 239
Among the remaining eight development tools projects, seven are Unity-based, which means they make use of considerable reusable code stored in serialized files. This results in a lower level of code cloning, as tools like NiCad are less effective at identifying clones within serialized files, which differ from standard source code. Project-7 (“lovr”), although written in C, also contains many model class files (/modules), which contributes to its relatively lower levels of code cloning. This observation does not contradict our conclusions.
In terms of Project-11 and Project-24, we find that they have relatively few serialized files in relation to their source code volume, causing higher levels of code cloning. In fact, Project-24, being an AR project, constructs models from real-world data rather than serialized files. Similarly, Project-11, an acoustic simulation and visualization project, does not require extensive use of serialized files. The remaining 18 projects maintain a relatively low level of code cloning.
In summary, the code cloning levels of VR software are generally lower than those of traditional software. This is because processes that tend to generate clones, such as class and method creation, are often implicitly stored within asset files. As a result, the cloning metrics might misleadingly appear lower. The impact of asset files and their relationship to cloning cannot be ignored, adding further challenges for clone detection in VR software.
Insight 2.  The main difference between code clones in VR software and regular software lies in the presence of serialized files, which significantly reduce the level of source code cloning in a project.
4.1.3 RQ3: How to identify the most appropriate metrics for VR cloning analysis?
To answer this RQ, we sequentially analyze the applicability of the metrics defined in Section 3.2.1.
$\sf NCF$ and $\sf NCC$. The physical interpretation of ${\sf NCF}$ is relatively intuitive; higher values directly indicate a greater cloning volume. However, the implication of ${\sf NCC}$ is less straightforward. Using Type3-2 in Table 4 as an example, we observe a notable difference between $\sf NCF$ and $\sf NCC$. For instance, comparing Project-2 to Project-1, the ${\sf NCF}$ ratio is approximately $2699/4448\approx 60.7\%$, whereas the ${\sf NCC}$ ratio is considerably lower at $326/1226\approx 26.6\%$. In contrast, when comparing Project-4 to Project-3, the ${\sf NCF}$ ratio is $409/462\approx 88.5\%$, but the ${\sf NCC}$ ratio is even higher at $163/166\approx 98.2\%$. To uncover the underlying cause of these differences, we analyze the correlation between the number of clone classes and their respective sizes. As shown in Table 7, Project-2 exhibits a significantly higher proportion of large-size (>100) and mid-size ([11, 100]) classes compared to Project-1, leading to a notable reduction in $\sf NCC$. However, Project-3 and Project-4 have similar clone class distributions across different intervals, which results in a close ${\sf NCC}$. The observations imply that ${\sf NCC}$ acts as a valuable indicator of clone class size, with lower values typically reflecting the presence of larger clone classes.
Table 7: The distribution of the sizes of clone classes.
Project Name NCC Interval 2 [3,10] [11, 100] 100 Sum
BlenderXR (PID:1) 898(73.2%) 304(24.8%) 23(1.9%) 1(0.1%) 1226
gpac (PID:2) 207(63.5%) 100(30.7%) 13(4.0%) 6(1.8%) 326
The-Seed-Link-Future (PID:3) 122(73.5%) 39(23.5%) 5(3.0%) 0 166
open-brush (PID:4) 123(75.5%) 37(22.7%) 3(1.8%) 0 163
Figure 2: The changes of ${\sf rcf}$ and ${\sf rcc}$ across diverse projects.
$\sf rcf$ and $\sf rcc$. To explore the implications of $\sf rcf$(i.e., the proportion of clone functions, representing the density of function-level code clones) and $\sf rcc$(i.e., the proportion of unique clone classes, representing the diversity of code clone patterns), we analyze their variations across different projects, as illustrated in Figure 2. The results reveal that ${\sf rcf}$ exhibits significant fluctuations, while ${\sf rcc}$ remains relatively consistent across projects except for Project-11. The anomaly in the project “SoundSpace” arises from two highly similar directories, “SoundSpace” and “rv_SoundSpace”, with 902 of 965 clone classes being pairwise clones. These predominantly size-2 clones lead to an unusually high ${\sf rcc}$. Additionally, no clear correlation is observed between ${\sf rcf}$ and ${\sf rcc}$, suggesting that the density of function-level code clones and the diversity of clone patterns are largely independent of each other. This decoupling highlights the complexity of code clone characteristics and underscores the need for separate consideration of density and diversity in clone analysis.
Figure 3: The changes of $\sf rca$, $\sf rcg$ and $\sf CI$ across diverse projects.
$\sf NCA$ and $\sf NCG$. The meanings of ${\sf NCA}$ and ${\sf NCG}$ are analogous to ${\sf NCF}$ and ${\sf NCC}$ in source code cloning. A higher ${\sf NCA}$ indicates a greater quantity of cloned assets within the project, while the impact of ${\sf NCG}$ is reflected in the concentration of clones. Given a fixed ${\sf NCA}$ value, a lower ${\sf NCG}$ suggests that the number of members in individual clone groups is higher, indicating more concentrated cloning within the project.
$\sf rca$, $\sf rcg$ and $\sf CI$. To explore the meanings of these three metrics, we plot their respective values for the 20 selected projects, as shown in Figure 3. The results indicate that ${\sf rca}$ and ${\sf CI}$ exhibit a certain degree of correlation, while ${\sf rcg}$ remains relatively stable across the 20 projects. We observe that Project-14 (“Situated-Empathy-in-VR”) and Project-25 (“Procrastination-VR”) show exceptionally high ${\sf CI}$, standing out significantly from other projects. Upon further investigation, we find that these two projects have the highest ${\sf rca}$ values (75.60% and 91.63%) among all projects, while their ${\sf rcg}$ values are the lowest (2.87% and 2.51%), meaning that they possess the most cloned files but the fewest clone groups. This corresponds to the case where a few files are heavily cloned, leading to the emergence of some exceptionally large clone groups. Thus, we can deduce that ${\sf CI}$ is jointly influenced by ${\sf rca}$ and ${\sf rcg}$, where ${\sf rca}$ shows a positive correlation and ${\sf rcg}$ a negative correlation.
Insight 3.  The proposed metrics measure cloning from distinct perspectives: the number or ratio of clone functions (files) measures the quantity of clones, while the number or ratio of clone classes (groups) evaluates clone concentration. By combining these metrics, a comprehensive clone evaluation can be achieved.
4.2 Source Code Cloning Analysis
4.2.1 RQ4: Which programming languages dominate VR software development and are more susceptible to code cloning?
To answer this RQ, we visualize the programming languages used by these software projects in Figure 4. The left pie chart shows the programming language distribution for all VR-345 projects, while the right highlights filtered high-star projects with star >= 200. In this context, the programming language mentioned for a project refers to the primary language used within the project.
(a) All VR-345 projects.
(b) 83 high-star projects.
Figure 4: Distribution of programming languages in VR projects.
We observe that C# dominates general VR projects, while it accounts for slightly over half in high-star projects. This is due to the fact that high-star projects not only feature engine-based applications but also include a substantial number of development tool projects, which are typically developed in various programming languages and serve as support for other projects. The advantages of C# language are most likely attributed to its close integration with the Unity game engine, which has emerged as one of the leading tools in VR development due to its ease of use and cross-platform features. Given that C# serves as the primary programming language for Unity, it is unsurprising that it has become the most widely adopted language for VR projects. Other languages employed in VR development include Python, C, Java, and Swift. In contrast, languages such as PHP, Ruby, WSDL, and ATL are absent. This is reasonable because web-oriented languages like PHP, Ruby, and WSDL are not well-suited for graphics-intensive applications due to limitations in execution efficiency and performance, while ATL is primarily designed for Windows-specific programming. It is also worth noting that JavaScript plays a significant role in VR project development. However, its absence in this analysis stems from a limitation in NiCad, which currently does not support clone detection for JavaScript.
To further examine which programming languages are more prone to code cloning, we identify projects with 10 or more clone classes ($\sf NCC$>= 10) in terms of Clone Type 3-2 within high-star VR projects and categorize these projects by their primary programming languages. Table 8 presents the distribution of projects across various languages at different cloning levels. The results reveal that C# accounts for the highest number of projects (18) with over 10 clone classes, followed by C with 5 projects. This finding aligns with the prominent role of C# as the most commonly used language in VR software development. However, when considering the cloning ratio, C ($5/11\approx 45.5\%$) surpasses C# ($18/46\approx 39.1\%$), indicating that C may have a higher likelihood of triggering cloning issues. This may stem from the fact that C, being process-oriented, does not support the code reuse mechanisms inherent to object-oriented languages, leading to more code cloning. In contrast, the presence of serialized files in C# programs leads to the occurrence of code clones at lower levels, as demonstrated in RQ2.
Insight 4.  C# is the most commonly used programming language in open-source VR projects, whereas C is more likely to cause source code cloning issues. However, C# introduces an additional challenge with the cloning of serialized files.
Table 8: The distribution of projects across different cloning intervals with regard to ${\sf NCC}$ using various languages.
Language NCC Interval [10, 30] [31, 50] [51, 100] 100 Sum
Java 1 0 0 0 1 (out of 10)
Python 3 0 0 1 4 (out of 14)
C 1 1 1 2 5 (out of 11)
C# 7 7 2 2 18 (out of 46)
4.2.2 RQ5: How do third-party libraries impact code cloning in VR software?
Table 9: The evaluation results of third-party libraries in the projects.
PID Project Library
[…]
11 SoundSpace 2225 965 SteamVR 70 3.17% 26 2.69%
12 RhythmAttack-VR 330 119 SteamVR 29 73.64% 8 68.07%
[…]
13 Dungeon-VR 254 98 SteamVR 29 57.87% 8 42.86%
[…]
15 mineRVa 197 71 SteamVR 33 97.46% 8 94.37%
[…]
16 VR-Escape-Room 120 46 SteamVR 29 91.67% 8 91.30%
[…]
17 Group6_ProjectNurture 248 82 Oculus 222 89.52% 69 84.15%
19 Terminal 160 49 SteamVR 22 56.88% 6 46.94%
[…]
20 elite-vr-cockpit 109 39 SteamVR 83 76.15% 28 71.79%
21 VRHamsterBall 74 28 SteamVR 66 89.19% 24 85.71%
22 OpendagVR2 45 16 SteamVR 29 64.44% 8 50.00%
23 epicslash 54 19 SteamVR 26 48.15% 7 36.84%
26 AirAttack 94 30 Oculus 54 57.45% 13 43.33%
To answer this RQ, we search through the 20 application projects to identify the presence of widely used third-party libraries (e.g., SteamVR, Oculus, OpenVR, Vive, Google VR, VRTK). The results are summarized in Table 9, listing only the projects with code clones induced by third-party libraries, along with the specific libraries involved.
Upon reviewing our results, we find that despite exploring various well-known VR third-party libraries, code clones are only successfully identified in projects using SteamVR, Oculus, and VRTK. Among these, SteamVR appears in 10 projects, while Oculus and VRTK each appear in 4 projects. However, when it comes to introducing source code clones, SteamVR has the lowest contribution, Oculus ranks in the middle, and VRTK contributes the most. This is reasonable because VRTK, as an open-source toolkit, supports numerous VR devices and platforms, providing high-level abstractions that simplify development and lead to significant code redundancy. On the other hand, SteamVR and Oculus are development kits designed for specific hardware, prioritizing compatibility and low-level optimization, which results in less cross-platform or high-level redundant code.
These findings confirm the presence of code clones caused by third-party libraries in VR software. Such clones are not introduced by the developers but are intrinsic to the third-party libraries themselves. For instance, in the three third-party libraries identified in our analysis, different projects display the same cloning results, such as (${\sf NCF}$:29, ${\sf NCC}$:8) of SteamVR in Project-12, 13, 16, 22, (${\sf NCF}$:54, ${\sf NCC}$:13) of Oculus in Project-12, 26, and (${\sf NCF}$:81, ${\sf NCC}$:34) of VRTK in Project-13, 16. These clones are triggered by the most commonly used code in these libraries, with additional clones being introduced by project-specific features. Although developers cannot modify clones in third-party tools, they can minimize them by using dependency injection, interface-based techniques, or opting for lightweight libraries when cross-platform support is unnecessary.
Insight 5.  Third-party libraries introduced in VR applications often contribute to source code cloning issues, sometimes constituting the majority of the cloning results. The more functionalities the third-party libraries provide, the higher the clone evaluation metrics tend to be.
4.2.3 RQ6: How do inter-version and intra-version code cloning evolve across different versions of a VR project?
To answer this RQ, we begin by examining whether code clones exist between different versions of the same project. To achieve this, we select two projects with a high number of versions and a wide range between versions for analysis. Specifically, We execute code clone detection on six sequential versions of each project, with the detection carried out between each pair of adjacent versions, yielding five data sets per project. The experimental results are illustrated in Figure 5.
(a) open-brush
(b) gpac
Figure 5: Detection results of inter-version code clones on five adjacent version pairs of two projects.
(a) gpac
(b) UnityPlugin
Figure 6: Detection results of intra-version clones across multiple versions when major version number changes.
(a) ViveInputUtility-Unity
(b) com.xrtk.core
Figure 7: Detection results of intra-version clones across multiple versions when secondary version number changes.
We find that in the “open-brush” project, the ${\sf rcc}$ and ${\sf rcf}$ values fall within the ranges of [0.158, 0.163] and [0.348, 0.359], respectively. In the “gpac” project, these values span wider ranges, with ${\sf rcc}$ between [0.168, 0.228] and ${\sf rcf}$ between [0.543, 0.664]. These observations suggest that while the degree of inter-version code cloning varies across projects, it remains consistently high, underscoring the prevalence of inter-version code clones in VR software development. This trend highlights that while code reuse improves development efficiency, it also poses the risk of propagating vulnerabilities, which may compromise software stability and security in future versions.
To examine intra-version code clones, we divide the analysis into three sub-RQs focusing on their evolution with changes in (i) major version numbers (e.g., 0.x.x, 1.x.x, 2.x.x), (ii) secondary version numbers (e.g., 0.1.x, 0.2.x, 0.3.x), and (iii) ending version numbers (e.g., 0.1.1, 0.1.2, 0.1.3). For each sub-RQ, we select two distinct projects and analyze their latest three eligible versions. These projects are chosen for their frequent release schedules, which provide a robust dataset for analysis.
For major version changes, Figure 6 and Figure 6 illustrate how ${\sf rcf}$ and ${\sf rcc}$ evolve with major version changes in the VR projects “gpac” and “UnityPlugin”, respectively. In the project “gpac”, ${\sf rcf}$ fluctuates by up to 0.036 and ${\sf rcc}$ by up to 0.005 across three major versions. Similarly, in the project “UnityPlugin”, ${\sf rcf}$ shows a variation of up to 0.007, and ${\sf rcc}$ up to 0.004. These observations suggest that the extent of code clones, as measured by ${\sf rcf}$ and ${\sf rcc}$, remains relatively stable during major version updates for VR projects.
For secondary version changes, Figure 7 and Figure 7 show the evolution of ${\sf rcf}$ and ${\sf rcc}$ as the secondary version numbers change in the projects “ViveInputUtility-Unity” and “com.xrtk.core”, respectively. We observe that in the project “ViveInputUtility-Unity”, ${\sf rcf}$ and ${\sf rcc}$ remain constant across the three analyzed versions. In the project “com.xrtk.core”, ${\sf rcf}$ fluctuates by up to 0.021, and ${\sf rcc}$ by up to 0.007. This indicates that the level of code clones generally remains stable, or even unchanged, during secondary version updates.
For ending version changes, we do not present figures because the variations in ${\sf rcf}$ and ${\sf rcc}$ across such updates are negligible. The minimal differences suggest that code clone patterns exhibit almost no significant evolution at this level of versioning.
Insight 6.   Inter-version code cloning is more common than intra-version code cloning, introducing significant security risks due to the potential propagation of vulnerabilities across multiple versions of the software.
4.3 Asset File Cloning Analysis
4.3.1 RQ7: How do cloning practices vary across different types of asset files in VR software?
To answer this RQ, we select four representative projects from distinct application categories and conduct clone detection on the serialized files within each project, with results illustrated in Figure 8.
We observe a considerable disparity between scene files and material files, with scene files typically exhibiting the lowest clone level, averaging ${\sf CI}\approx 0.1125$, while material files demonstrate the highest, with an average of ${\sf CI}\approx 1.05$. This is possibly due to the fact that VR scenes generally contain unique content and structures that are difficult to replicate or share. Conversely, material files are extensively reused across various objects and scenes for optimization and performance reasons.
Additionally, we find that the clone level of prefab files is intermediate, with an average of ${\sf CI}\approx 0.375$, yet it shows substantial fluctuations. This may be attributed to variations in system design and resource reuse strategies across different projects, which are heavily influenced by the developer’s style. Some projects may require complex, customized prefabs, while others rely on a large number of standardized prefabs that are heavily reused. The use of techniques like prefab variants and runtime instantiation in complex prefabs, analogous to object-oriented programming, can lead to a reduced ${\sf CI}$. In contrast, extensive reuse of standardized prefabs results in a higher ${\sf CI}$, facilitating development but introducing redundancy, as changes to attributes need to be applied to every file within the clone group.
Figure 8: Analysis of cloning behavior of various asset files from four typical projects.
Insight 7.  In VR software, the cloning behavior of asset files varies with complexity. Complex files like scenes have lower clone levels, simpler ones like materials have higher, and prefab files fluctuate with development style.
5 Discussion
5.1 Recommendations
5.1.1 For Researchers
We recommend that researchers should give thorough consideration to the distinctions between different software types and general software when conducting studies on code cloning, ensuring a more comprehensive exploration of these differences. In particular, VR software necessitates deeper exploration of the cloning phenomena in its specialized asset and configuration files. The current study utilizes large language model technology to examine cloning behaviors in these files, hoping to provide significant insights for future investigations in this domain.
5.1.2 For VR Software Developers
We recommend that VR software developers carefully address the issue of code cloning throughout the development process to mitigate potential software vulnerabilities and minimize maintenance challenges. This includes using design patterns and coding paradigms that are suitable for the project, adopting modular design, avoiding direct copying and pasting of code blocks or files, and utilizing static analysis tools for cloning evaluation during the software development process. Specifically, for the unique asset files in VR software, developers should also consider applying similar design patterns and coding paradigms, such as using variants or instantiations, to avoid direct duplication of asset files. Since these files cannot be detected by traditional detection tools, the large model detection method used in this study could be applied. By employing these strategies, developers can effectively prevent or mitigate code cloning issues, enhancing the software’s maintainability, scalability, and overall quality.
5.1.3 For Third-party Library Designers
VR software often includes cloned function classes within third-party libraries, which is particularly common in libraries with cross-platform capabilities. We recommend using specific techniques to mitigate these cloning phenomena. For example, platform abstraction layers or interface abstract classes can be used to encapsulate specific implementations, conditional compilation directives can compile different code blocks for different platforms, dependency injection, and factory classes can enable dynamic implementation, or cross-platform frameworks (e.g., OpenXR) can be employed. Additionally, projects should select third-party libraries that best meet their needs; for example, if the project only needs to run on a specific hardware platform, a lightweight third-party library suitable for that platform can be chosen.
5.2 Future Work
Future research on the topic of this study could be approached from several directions. First, regarding the asset file types in VR software, our research does not cover all file types. Some files, such as audio, images, textures, and others stored in specialized formats, are not thoroughly investigated. This could be a key direction for future studies. Secondly, our evaluation of source code and serialized asset file cloning in VR software is conducted separately. A possible future direction would be to integrate these two aspects into a unified system for a comprehensive assessment. Finally, although we use the best available cloning detection tools for source code, with the advancement of AI technologies, future research could explore the use of large language models for detecting both source code and serialized asset files together. A potential solution could involve designing a large language model agent specifically for cloning behaviors in VR software, which could also be integrated with the aforementioned comprehensive evaluation of both file types.
6 Related Work
6.1 Empirical Study of VR Software
The empirical study of VR software has gained increased attention in recent years, and the scope includes educational [71], marketing communication [72], user operational performance [73], automated testing [74], vulnerabilities [65, 75, 76] and code cloning [17]. Among the studies, Our prior work [17] provides the most relevant research on code cloning in open-source VR software. However, the study does not address the code cloning of serialized files, which represents an important aspect of VR software.
6.2 Code Clone Detection
Traditional methods for detecting code clones include text-based, token-based, and AST-based techniques. Tools like Simian [27], CCfinder [77, 13], and NiCad [18, 15] focus on comparing code at the syntactic or token level, while others like CloneDR [33, 34] and Duplication Finder [78, 79] leverage the Abstract Syntax Tree for deeper analysis of code structure.
Recent advancements have introduced the use of Large Language Models (LLMs), such as Codex [80, 81] and CodeBERT [41, 42], to detect clones with semantic awareness. These models understand code functionality, allowing them to identify clones that may not be syntactically identical but are semantically equivalent. By leveraging LLMs, clone detection can move beyond superficial matches, improving accuracy in detecting complex clones and offering greater flexibility across different programming languages. Recent research [45] shows that NiCad performs best in non-LLMs-based detection, while GPT-4 performs best in LLMs-based detection. Both of them achieve the highest precision and recall for each clone type in the same type of comparison.
6.3 File Clone Detection
Asset files in VR software, representing the differences from traditional software, can be considered specialized text files. There are several methods to detect their similarity, with some of the more representative ones being content-based methods, fingerprint-based methods, and semantic (deep learning) methods.
Content-based detection methods [48, 49, 50, 51] focus on directly comparing the file contents using traditional techniques such as string matching and text block comparison. However, these methods tend to be less efficient when dealing with large-scale systems. Fingerprint-based detection methods [53, 54] improve efficiency by generating fingerprints for the files and using efficient lookup algorithms, making them particularly suitable for large-scale source code repositories. However, they cannot detect clones caused by syntax or structural changes. Deep learning-based detection methods [56, 57, 58, 59] rely on neural networks to model the semantics of the code or text. These methods can identify clones that differ in structure but are semantically similar, offering strong adaptability and high accuracy. However, they come with greater implementation complexity and computational overhead. Nevertheless, with the recent development of LLM technology and architecture, the disadvantages of deep learning-based detection methods have been alleviated to a certain extent, making it an advantageous method for similarity detection of serialized text files.
7 Conclusion
In this paper, we conduct a quantitative empirical study on code cloning in open-source VR projects. We refine the VR-345 dataset from previous studies for this research, ensuring a balance between the influence and range of the projects. We then propose a set of metrics to measure code cloning from different perspectives. Finally, we perform empirical experiments on seven carefully designed research questions at three distinct levels, in order to capture an overview of code cloning in VR software.
Our study reveals the key differences between VR software and traditional software in terms of code cloning. Due to these differences, certain files cannot be detected by traditional detection tools, which is why we employ large language model technology to assist in our detection. We evaluate and analyze the experimental results from various aspects such as clone distribution, clone concentration, programming languages, third-party libraries, and the impact of different types of assets. Through this, we answer the seven proposed research questions, deepening our understanding of code cloning in VR software.